DIPOLE MOMENTS OF AMINE HYDROGEN HALIDE COMPLEXES

\underline{C. S. BRAUER}, M. B. CRADDOCK, E. M. GRUMSTRUP, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455; M. C. ORILALL, Department of Chemistry, Concordia University, Moorhead, MN 56562.

Amine hydrogen halide complexes are prototypical systems in which to study the transfer of a proton across a hydrogen bond. We report the results of Stark effect experiments using a pulsed nozzle Fourier transform microwave spectrometer to determine the dipole moments of: H_{3}^{1}^{5}N-H^{3}^{5}Cl (4.05865 ±0.00095 D), (CH_{3})_{3}^{1}^{5}N-H^{3}^{5}Cl (7.129 ±0.014 D), H_{3}^{1}^{5}N-H^{7}^{9}Br (4.2577 ±0.0022 D), and (CH_{3})_{3}^{1}^{5}N-H^{7}^{9}Br (8.396 ±0.013 D). The results will be discussed in terms of the degree of proton transfer, as previously elucidated by Legon and coworkers^{a} and should be of interest in view of a large body of matrix isolation work in which the interaction between the complex and the host matrix has been investigated.

aA. C. Legon Chem. Soc. Rev. \underline{22}(3), 153, (1993).